OpenStudy (anonymous):

How do I graph this? three factors: (3x + 1) (x -2) (x + 4). multiply them out to get a cubic equation ; and then you can use factorisation to get to the solution. what are the factors?

OpenStudy (anonymous):

now multiply them out to get the cubic equation

OpenStudy (anonymous):

and the factors are already given which are (3x + 1) (x -2) (x + 4).

OpenStudy (anonymous):

i didnt get exactly the question more explaination plz

OpenStudy (anonymous):

factors are 3x + 1 = 0; x = -1/3 x-2 = 0;x = 2 x+ 4 = 0;x=-4

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

I'm told to graph this...how do I do that?

OpenStudy (anonymous):

You can graph them by moving a curve thru all these three points on the the x axis...wait im drawing and sending a word file

OpenStudy (anonymous):

can you draw it? still confused:(

OpenStudy (anonymous):

I plot the -1/3, 2 ,+and -4 and then do the curve?

OpenStudy (anonymous):

-1/3, 2 and -4*

OpenStudy (anonymous):

thank you dear!!!!:))))

OpenStudy (anonymous):

heu look at it ..we will draw this graph like that

OpenStudy (anonymous):

let me see:)

OpenStudy (anonymous):

omg thank you thank you thank you thank you!!!!!!!!!!!!!!!!!!!!!!!!!! @ritez

OpenStudy (anonymous):

You are most welcome...Cheers!